OpenStudy (cherie_magee):

Which of the following is a solution of y - x < -3? (6, 2) (2, 6) (2, -1)

OpenStudy (cherie_magee):

@FriedRice can you help me with this????

OpenStudy (anonymous):

well add x to both sides to get y<-3+x then use the plug in x and y to find which one comes out equal

OpenStudy (anonymous):

so -3+6=3 and -3+2=-1 and since you are looking for y<-3+x your answer would be (2,6)

OpenStudy (anonymous):

sorry i meant (2,-1)

OpenStudy (anonymous):

well you have the equation y - x < -3 right? it mostly saying y - x = -3 and then you can say y =x - 3 and just plug in the point and see what fits the best

OpenStudy (anonymous):

so Kaylei was right it is (2,-1)